### v1

@@ 3 Posterior smoothing with dimension-independent error
$W_{1}(\mu,\nu):=\inf_{\gamma\in\Pi(\mu,\nu)}\int\left\|\mathbf{x}-\mathbf{y}\right\|\gamma(\mathbf{x},\mathbf{y})\mathrm{d}\mathbf{x}\mathrm{d}\mathbf{y},$
@@ 4 Dimension-independent block acceptance and convergence rate
It is shown in [32] that MLwG given in Algorithm 1 has dimension-independent block acceptance rate and dimension-independent convergence rate for smooth densities under assumptions on sparse conditional structure and blockwise log-concavity. However, due to the TV prior, the extension to the smoothed posterior density $\pi_{\varepsilon}$ in Eq. 14 is not straight-forward. In the following, we present two results which show that we can obtain similar dimension-independent results for $\pi_{\varepsilon}$ by using the concept of $c$-diagonal block dominance in 1.
@@
Here $M$ is a dimension-independent constant depending on the block size $q$, $C_{\mathbf{A}},\delta,\varepsilon$, and $\max_{j}\left\|[\mathbf{A}\mathbf{x}^{k,i-1}-\mathbf{y}]_{j}\right\|$.
Proof 4.2.
@@ 5 Local & parallel MLwG algorithm
The posterior Eq. 14 can be interpreted a Gibbs density over an undirected graph. The nodes are given by the pixels, and the edges are given by the neighborhoods of the pixels. In particular, a pixel is connected to all its neighbors. Pixels are neighbors to each other if their associated random variables directly interact with each other in the joint density. For example, two random variables $Z_{1}\in\mathbb{R}$ and $Z_{2}\in\mathbb{R}$ directly interact in the density $\rho(Z)\propto\exp(-Z_{1}Z_{2})$ but not in $\rho(Z)\propto\exp(-(Z_{1}+Z_{2}))$. A set of pixels in which all pixels are neighbors to each other is called clique. Furthermore, a maximal clique is a clique to which no pixel can be added such that the set is still a clique [18].
@@ 5.1 Local block likelihood
${\mathbf{U}_{i}^{+r}}^{\rm T}\mathbf{x}_{i}^{+rr}$ $=\mathbf{x}_{i}^{+r}$
${\mathbf{U}_{i}}^{\rm T}\mathbf{x}_{i}^{+rr}$ $=\mathbf{x}_{i}.$
@@ 5.2 Local block prior
The set of pixels $\Phi_{i}$ on which the conditional Eq. 20 depends is determined by the convolution. To compute the TV in $\varphi_{\varepsilon}^{k,i-1}$, we do not require all pixels in $\Phi_{i}$, but only the pixels of block $i$, extended by a frame of one pixel around it. In the following, we denote this extended block by $\mathbf{x}_{i}^{+1}\in\mathbb{R}^{q_{i}^{+1}}$. Note that the extended block $\mathbf{x}_{i}^{+1}$ can have different sizes $q_{i}^{+1}$, depending on whether it is located in a corner, at an edge or in the interior of the image. Moreover, we define the component selection matrix $\mathbf{W}_{i}\in\mathbb{R}^{q_{i}^{+1}\times q_{i}}$, which allows for the following mapping between $\mathbf{x}_{i}$ and $\mathbf{x}_{i}^{+1}$:
@@
We show an illustration of the extended block in Figure 3.
Figure 3: Examples of extended blocks for an interior block (left) and a corner block (right) in order to compute the local block prior.
@@
$\mathbf{D}_{i}^{l}$ $={\mathbf{D}_{i}^{l}}^{+1}\mathbf{W}_{i}$
$\mathbf{b}_{l}^{k,i-1}$ $={\mathbf{D}_{i}^{l}}^{+1}\mathbf{P}_{i}^{+1}\mathbf{x}_{i}^{+1}$
@@ 5.3 Local & parallel MLwG algorithm
6:           Get $\mathbf{x}_{i}^{+rr}$ and $\mathbf{x}_{i}^{+1}$ from the current state $\mathbf{x}^{k,i}$ (see Figs. 2 and 3 for the examples).
@@ Appendix A Proofs
A.1 Proof of Lemma 5.1
A.2 dimension-independent approximation error
@@
$\sum_{k}\left\|\nabla_{k}u\right\|_{L^{\infty}}\leq 2\lambda^{-1}c^{-1}.$
@@
$\int\phi(\mathbf{x}_{i})\pi(\mathbf{x})\mathrm{d}\mathbf{x}=~{}$ $\int\left(\Delta u(\mathbf{x})+\nabla\log\pi_{\varepsilon}(\mathbf{x})\cdot\nabla u(\mathbf{x})\right)\pi(\mathbf{x})\mathrm{d}\mathbf{x}$
@@
[…]
Recall the definition of the Wasserstein-$1$ distance, we have
(33) $\begin{split}\max_{i}W_{1}(\pi_{i},\pi_{\varepsilon,i})=~{}&\max_{i}\sup_{\phi_{0}\in{\rm Lip}_{1}(\mathbb{R}^{q})}\int\phi_{0}(\mathbf{x}_{i})\left(\pi_{i}(\mathbf{x}_{i})-\pi_{\varepsilon,i}(\mathbf{x}_{i})\right)\mathrm{d}\mathbf{x}_{i}\\ =~{}&\max_{i}\sup_{\phi\in{\rm Lip}_{1}^{0}(\pi_{\varepsilon,i})}\int\phi(\mathbf{x}_{i})\pi(\mathbf{x}_{i},\mathbf{x}_{\backslash i})\mathrm{d}\mathbf{x}_{i}\mathrm{d}\mathbf{x}_{\backslash i}\\ \leq~{}&2\lambda^{-1}c^{-1}\max_{j}\left\|\nabla_{j}\log\pi_{\varepsilon}-\nabla_{j}\log\pi\right\|_{L^{1}(\pi)}.\end{split}$
@@
$\qquad\cdot\left(\left(\mathbf{D}_{\alpha}^{(v)}\mathbf{x}\right)\left(\mathbf{D}_{\alpha,j}^{(v)}\right)^{\rm T}+\left(\mathbf{D}_{\alpha}^{(h)}\mathbf{x}\right)\left(\mathbf{D}_{\alpha,j}^{(h)}\right)^{\rm T}\right)$
[…]
Here we denote $\mathbf{D}_{\alpha,j}^{(v)}\in\mathbb{R}^{1\times q}$ as the $j$-th block of $\mathbf{D}_{\alpha}\in\mathbb{R}^{1\times d}$, and similar for $\mathbf{D}_{\alpha,j}^{(h)}$. To control $\left\|\mathbf{v}_{\alpha}(\mathbf{x})\right\|$, notice
$\left\|\left(\mathbf{D}_{\alpha}^{(v)}\mathbf{x}\right)\left(\mathbf{D}_{\alpha,j}^{(v)}\right)^{\rm T}+\left(\mathbf{D}_{\alpha}^{(h)}\mathbf{x}\right)\left(\mathbf{D}_{\alpha,j}^{(h)}\right)^{\rm T}\right\|$
[…]
@@
(34) $\begin{split}\|\nabla_{j}\log\pi(\mathbf{x})-\nabla_{j}\log&\pi_{\varepsilon}(\mathbf{x})\|_{L^{1}(\pi)}\leq\delta\cdot\mathbb{E}_{\mathbf{x}\sim\pi}\sum_{\alpha:\mathbf{v}_{\alpha}(\mathbf{x})\neq 0}\left\|\mathbf{v}_{\alpha}(\mathbf{x})\right\|\\ \leq~{}&2\delta\cdot\mathbb{E}_{\mathbf{x}\sim\pi}\sum_{\alpha:\mathbf{v}_{\alpha}(\mathbf{x})\neq 0}\left[1-\frac{\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|}{\sqrt{\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|^{2}+\varepsilon}}\right]\\ \leq~{}&2\delta(m+2)^{2}\max_{\alpha}\mathbb{E}_{\mathbf{x}\sim\pi}\left[1-\frac{\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|}{\sqrt{\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|^{2}+\varepsilon}}\right].\end{split}$
Denote the function
$I_{\alpha}(t):=\mathbb{E}_{\mathbf{x}\sim\pi}\left[1-\frac{\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|}{\sqrt{\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|^{2}+t^{2}}}\right].$
Then $I_{\alpha}(0)=0$ and
$I_{\alpha}^{\prime}(t)=~{}$ $\mathbb{E}_{\mathbf{x}\sim\pi}\left[\frac{\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|t}{\left(\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|^{2}+t^{2}\right)^{3/2}}\right]\leq\mathbb{E}_{\mathbf{x}\sim\pi}\left[\frac{\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|^{2}+t^{2}}{2\left(\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|^{2}+t^{2}\right)^{3/2}}\right]$
@@
$\max_{\alpha}\mathbb{E}_{\mathbf{x}\sim\pi}\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|^{-1}\leq C_{\pi}.$
@@
$\mathbb{E}_{\mathbf{x}\sim\pi}\left[1-\frac{\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|}{\sqrt{\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|^{2}+\varepsilon}}\right]\leq\frac{1}{2}C_{\pi}\varepsilon^{1/2}.$
@@
A.2.2 Some technical lemmas
@@
(35) $\sum_{k}\left\|\nabla_{k}u\right\|_{L^{\infty}}\leq 2\lambda^{-1}c^{-1}.$
@@
It suffices to prove Eq. 35 for $\phi\in C^{1}\cap{\rm Lip}_{1}^{0}(\pi_{\varepsilon,i})$. Note this space is dense in ${\rm Lip}_{1}^{0}(\pi_{\varepsilon,i})$, so for general $\phi\in{\rm Lip}_{1}(\pi_{\varepsilon,i})$, we can take a sequence of $\phi_{k}\in C^{1}\cap{\rm Lip}_{1}^{0}(\pi_{\varepsilon,i})$ that converges to $\phi$. Eq. 35 holds uniformly for $\phi_{k}$, so that passing the limit shows that it holds for any $\phi\in{\rm Lip}_{1}^{0}(\pi_{\varepsilon,i})$.
Now fix any $\phi\in C^{1}\cap{\rm Lip}_{1}^{0}(\pi_{\varepsilon,i})$. It is straightforward to verify by standard elliptic theory that the solution exists (up to a constant) and $u\in C^{3}$. Taking gradient in Eq. 32,
$\nabla\phi(\mathbf{x}_{i})=~{}$ $\Delta\left(\nabla u(\mathbf{x})\right)+\nabla^{2}\log\pi_{\varepsilon}(\mathbf{x})\cdot\nabla u(\mathbf{x})+\nabla^{2}u(\mathbf{x})\cdot\nabla\log\pi_{\varepsilon}(\mathbf{x})$
[…]
The $j$-th block reads
@@
Recall by definition, $\pi_{\varepsilon}(\mathbf{x})\propto\exp(-l(\mathbf{x})-\varphi_{\varepsilon}(\mathbf{x}))$, where
$l(\mathbf{x})=\frac{\lambda}{2}\left\|\mathbf{y}-\mathbf{A}\mathbf{x}\right\|_{2}^{2},\quad\varphi_{\varepsilon}(\mathbf{x})=\delta\sum_{\alpha=1}^{d}\sqrt{(\mathbf{D}_{\alpha}^{(v)}\mathbf{x})^{2}+(\mathbf{D}_{\alpha}^{(h)}\mathbf{x})^{2}+\varepsilon}.$
Direct computation shows
@@
For simplicity, denote $H(\mathbf{x})=\nabla^{2}\varphi_{\varepsilon}(\mathbf{x})$. Then Eq. 36 reads
$\nabla_{j}\phi(\mathbf{x}_{i})=\mathcal{L}(\nabla_{j}u)(\mathbf{x})-\lambda\sum_{k}(\mathbf{A}^{\rm T}\mathbf{A})_{jk}\nabla_{k}u(\mathbf{x})-\sum_{k}H_{jk}\nabla_{k}u(\mathbf{x}).$
Multiple from left by $-\left(\nabla_{j}u(\mathbf{x})\right)^{\rm T}$,
(38) $\begin{split}&-\left(\nabla_{j}u(\mathbf{x})\right)^{\rm T}\nabla_{j}\phi(\mathbf{x}_{i})+\left(\nabla_{j}u(\mathbf{x})\right)^{\rm T}\mathcal{L}(\nabla_{j}u)(\mathbf{x})\\ =~{}&\lambda\sum_{k}\left(\nabla_{j}u(\mathbf{x})\right)^{\rm T}(\mathbf{A}^{\rm T}\mathbf{A})_{jk}\nabla_{k}u(\mathbf{x})+\sum_{k}\left(\nabla_{j}u(\mathbf{x})\right)^{\rm T}H_{jk}\nabla_{k}u(\mathbf{x}).\end{split}$
Since $\mathbf{A}^{\rm T}\mathbf{A}$ is $c$-diagonal block dominant with matrix $M\in\mathbb{R}^{b\times b}$, it holds that
$\sum_{k}\left(\nabla_{j}u(\mathbf{x})\right)^{\rm T}(\mathbf{A}^{\rm T}\mathbf{A})_{jk}\nabla_{k}u(\mathbf{x})$
@@
For the second in the right hand side of Eq. 38, notice $H$ is positive definite since $\varphi_{\varepsilon}$ is convex, so that $H_{jj}$ is also positive definite. Therefore,
$\sum_{k}\left(\nabla_{j}u(\mathbf{x})\right)^{\rm T}H_{jk}(\mathbf{x})\nabla_{k}u(\mathbf{x})$

### v2

@@ 3 Posterior smoothing with dimension-independent error
$W_{1}(\mu,\nu):=\inf_{\gamma\in\Pi(\mu,\nu)}\int\left\|\mathbf{x}-\mathbf{y}\right\|_{2}\gamma(\mathbf{x},\mathbf{y})\mathrm{d}\mathbf{x}\mathrm{d}\mathbf{y},$
@@ 4 Dimension-independent block acceptance and convergence rate
It is shown in [32] that MLwG given in Algorithm 1 has dimension-independent block acceptance rate and dimension-independent convergence rate for smooth densities under assumptions on sparse conditional structure and blockwise log-concavity. In the following, we present two results which show that we can obtain similar dimension-independent results for $\pi_{\varepsilon}$ by using the concept of $c$-diagonal block dominance in 1. Since the proofs are rather similar as the ones for Proposition 3.3 and Theorem 3.6 in [32], we skip them here.
@@
Here $M$ is a dimension-independent constant depending on the block size $q$, $C_{\mathbf{A}},\delta,\varepsilon$, and $\max_{j}\left\|[\mathbf{A}\mathbf{x}^{k,i-1}-\mathbf{y}]_{j}\right\|_{2}$.
@@ 5 Local & parallel MLwG algorithm
The posterior Eq. 14 can be interpreted as a Gibbs density over an undirected graph. The nodes are given by the pixels, and the edges are given by the neighborhoods of the pixels. In particular, a pixel is connected to all its neighbors. Pixels are neighbors to each other if their associated random variables directly interact with each other in the joint density. For example, two random variables $Z_{1}\in\mathbb{R}$ and $Z_{2}\in\mathbb{R}$ directly interact in the density $\rho(Z)\propto\exp(-Z_{1}Z_{2})$ but not in $\rho(Z)\propto\exp(-(Z_{1}+Z_{2}))$. A set of pixels in which all pixels are neighbors to each other is called clique. Furthermore, a maximal clique is a clique to which no pixel can be added such that the set is still a clique [18].
@@ 5.1 Local block likelihood
${\mathbf{U}_{i}^{+r}}^{\rm T}\mathbf{x}_{i}^{+rr}=\mathbf{x}_{i}^{+r},\qquad{\mathbf{U}_{i}}^{\rm T}\mathbf{x}_{i}^{+rr}=\mathbf{x}_{i}.$
@@ 5.2 Local block prior
The set of pixels $\Phi_{i}$ on which the conditional Eq. 20 depends is determined by the convolution. To compute the TV in $\varphi_{\varepsilon}^{k,i-1}$, we do not require all pixels in $\Phi_{i}$, but only the pixels of block $i$, extended by a frame of one pixel around it in the similar way as illustrated in Figure 2 with $12r=$ 1. In the following, we denote this extended block by $\mathbf{x}_{i}^{+1}\in\mathbb{R}^{q_{i}^{+1}}$. Note that the extended block $\mathbf{x}_{i}^{+1}$ can have different sizes $q_{i}^{+1}$, depending on whether it is located in a corner, at an edge or in the interior of the image. Moreover, we define the component selection matrix $\mathbf{W}_{i}\in\mathbb{R}^{q_{i}^{+1}\times q_{i}}$, which allows for the following mapping between $\mathbf{x}_{i}$ and $\mathbf{x}_{i}^{+1}$:
@@
$\mathbf{D}_{i}^{l}={\mathbf{D}_{i}^{l}}^{+1}\mathbf{W}_{i},\qquad\mathbf{b}_{l}^{k,i-1}={\mathbf{D}_{i}^{l}}^{+1}\mathbf{P}_{i}^{+1}\mathbf{x}_{i}^{+1},$
@@ 5.3 Local & parallel MLwG algorithm
6:           Get $\mathbf{x}_{i}^{+rr}$ and $\mathbf{x}_{i}^{+1}$ from the current state $\mathbf{x}^{k,i}$ (see Fig. 2 for example).
@@ Appendix A Proofs
A.1 Dimension-independent approximation error
@@
$\sum_{j}\left\|\nabla_{j}u\right\|_{L^{\infty}}\leq 2\lambda^{-1}c^{-1}.$
@@
$\int\phi(\mathbf{x}_{i})\pi(\mathbf{x})\mathrm{d}\mathbf{x}=~{}$ $\int\left(-\nabla u(\mathbf{x})\cdot\nabla\log\pi(\mathbf{x})+\nabla\log\pi_{\varepsilon}(\mathbf{x})\cdot\nabla u(\mathbf{x})\right)\pi(\mathbf{x})\mathrm{d}\mathbf{x}$
@@
Combine the above two inequalities, we have
(32) $\begin{split}\max_{i}W_{1}(\pi_{i},\pi_{\varepsilon,i})=~{}&\max_{i}\sup_{\phi\in{\rm Lip}_{1}^{0}(\pi_{\varepsilon,i})}\int\phi(\mathbf{x}_{i})\pi(\mathbf{x})\mathrm{d}\mathbf{x}\\ \leq~{}&2\lambda^{-1}c^{-1}\max_{j}\left\|\nabla_{j}\log\pi_{\varepsilon}-\nabla_{j}\log\pi\right\|_{L^{1}(\pi)}.\end{split}$
@@
$\qquad\cdot\left(\left(\mathbf{D}_{\alpha}^{(v)}\mathbf{x}\right)\left(\mathbf{D}_{\alpha,j}^{(v)}\right)^{\rm T}+\left(\mathbf{D}_{\alpha}^{(h)}\mathbf{x}\right)\left(\mathbf{D}_{\alpha,j}^{(h)}\right)^{\rm T}\right)=:\delta\sum_{\alpha=1}^{d}\mathbf{v}_{\alpha}(\mathbf{x}).$
Here we denote $\mathbf{D}_{\alpha,j}^{(v)}\in\mathbb{R}^{1\times q}$ as the $j$-th block of $\mathbf{D}_{\alpha}\in\mathbb{R}^{1\times d}$, and similar for $\mathbf{D}_{\alpha,j}^{(h)}$. To control $\left\|\mathbf{v}_{\alpha}(\mathbf{x})\right\|_{2}$, notice
$\left\|\left(\mathbf{D}_{\alpha}^{(v)}\mathbf{x}\right)\left(\mathbf{D}_{\alpha,j}^{(v)}\right)^{\rm T}+\left(\mathbf{D}_{\alpha}^{(h)}\mathbf{x}\right)\left(\mathbf{D}_{\alpha,j}^{(h)}\right)^{\rm T}\right\|_{2}\leq|\mathbf{D}_{\alpha}^{(v)}\mathbf{x}|\|\mathbf{D}_{\alpha,j}^{(v)}\|_{2}+|\mathbf{D}_{\alpha}^{(h)}\mathbf{x}|\|\mathbf{D}_{\alpha,j}^{(h)}\|_{2}$
$\leq\sqrt{2}\left(|\mathbf{D}_{\alpha}^{(v)}\mathbf{x}|+|\mathbf{D}_{\alpha}^{(h)}\mathbf{x}|\right)\leq 2\left((\mathbf{D}_{\alpha}^{(v)}\mathbf{x})^{2}+(\mathbf{D}_{\alpha}^{(h)}\mathbf{x})^{2}\right)^{1/2}=:2\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|_{2}$
@@
(33) $\begin{split}\|\nabla_{j}\log\pi(\mathbf{x})-\nabla_{j}\log&\pi_{\varepsilon}(\mathbf{x})\|_{L^{1}(\pi)}\leq\delta\cdot\mathbb{E}_{\mathbf{x}\sim\pi}\sum_{\alpha:\mathbf{v}_{\alpha}(\mathbf{x})\neq 0}\left\|\mathbf{v}_{\alpha}(\mathbf{x})\right\|_{2}\\ \leq~{}&2\delta(m+2)^{2}\max_{\alpha}\mathbb{E}_{\mathbf{x}\sim\pi}\left[1-\frac{\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|_{2}}{\sqrt{\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|_{2}^{2}+\varepsilon}}\right].\end{split}$
Denote the function $I_{\alpha}(t):=\mathbb{E}_{\mathbf{x}\sim\pi}\left[1-\frac{\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|_{2}}{\sqrt{\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|_{2}^{2}+t^{2}}}\right]$. Then $I_{\alpha}(0)=0$ and
$I_{\alpha}^{\prime}(t)=~{}$ $\mathbb{E}_{\mathbf{x}\sim\pi}\left[\frac{\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|_{2}t}{\left(\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|_{2}^{2}+t^{2}\right)^{3/2}}\right]\leq\mathbb{E}_{\mathbf{x}\sim\pi}\left[\frac{\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|_{2}^{2}+t^{2}}{2\left(\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|_{2}^{2}+t^{2}\right)^{3/2}}\right]$
@@
$\max_{\alpha}\mathbb{E}_{\mathbf{x}\sim\pi}\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|_{2}^{-1}\leq C_{\pi}.$
@@
$\mathbb{E}_{\mathbf{x}\sim\pi}\left[1-\frac{\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|_{2}}{\sqrt{\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|_{2}^{2}+\varepsilon}}\right]\leq\frac{1}{2}C_{\pi}\varepsilon^{1/2}.$
@@
A.1.2 Some technical lemmas
@@
(34) $\sum_{j}\left\|\nabla_{j}u\right\|_{L^{\infty}}\leq 2\lambda^{-1}c^{-1}.$
@@
It suffices to prove Eq. 34 for $\phi\in C^{1}\cap{\rm Lip}_{1}^{0}(\pi_{\varepsilon,i})$. Note this space is dense in ${\rm Lip}_{1}^{0}(\pi_{\varepsilon,i})$, so for general $\phi\in{\rm Lip}_{1}(\pi_{\varepsilon,i})$, we can take a sequence of $\phi_{k}\in C^{1}\cap{\rm Lip}_{1}^{0}(\pi_{\varepsilon,i})$ that converges to $\phi$. Eq. 34 holds uniformly for $\phi_{k}$, so that passing the limit shows that it holds for any $\phi\in{\rm Lip}_{1}^{0}(\pi_{\varepsilon,i})$. Now fix any $\phi\in C^{1}\cap{\rm Lip}_{1}^{0}(\pi_{\varepsilon,i})$. It is straightforward to verify by standard elliptic theory that the solution exists (up to a constant) and $u\in C^{3}$. Taking gradient w.r.t. $\mathbf{x}_{j}$ in Eq. 31, we obtain
@@
Recall by definition Eq. 14, $\pi_{\varepsilon}(\mathbf{x})\propto\exp(-l(\mathbf{x})-\varphi_{\varepsilon}(\mathbf{x}))$. Direct computation shows
@@
For simplicity, denote $H(\mathbf{x})=\nabla^{2}\varphi_{\varepsilon}(\mathbf{x})$. Multiplying Eq. 35 from left by $-\left(\nabla_{j}u(\mathbf{x})\right)^{\rm T}$,
(37) $\begin{split}-\left(\nabla_{j}u(\mathbf{x})\right)^{\rm T}\nabla_{j}\phi(\mathbf{x}_{i})+&\left(\nabla_{j}u(\mathbf{x})\right)^{\rm T}\mathcal{L}(\nabla_{j}u)(\mathbf{x})\\ =~{}&\sum_{k}\left(\nabla_{j}u(\mathbf{x})\right)^{\rm T}(\lambda\mathbf{A}^{\rm T}\mathbf{A}+H(\mathbf{x}))_{jk}\nabla_{k}u(\mathbf{x}).\end{split}$
Since $\mathbf{A}^{\rm T}\mathbf{A}$ is $c$-diagonal block dominant with matrix $M\in\mathbb{R}^{b\times b}$, and $H_{jj}$ is positive definite since $\varphi_{\varepsilon}$ is convex, so that,
$\sum_{k}\left(\nabla_{j}u(\mathbf{x})\right)^{\rm T}(\lambda\mathbf{A}^{\rm T}\mathbf{A}+H(\mathbf{x}))_{jk}\nabla_{k}u(\mathbf{x})$
@@
Next we control $\|H_{jk}(\mathbf{x})\|_{2}$. Note $\mathbf{D}^{(v)}_{\alpha,i}=0$ if $\alpha$ does not live in block $i$ or a vertical neighbor of block $i$. Similarly for $\mathbf{D}^{(h)}_{\alpha,i}$. Therefore, $H_{ij}(\mathbf{x})=0$ if $i$ and $j$ are not neighbors. For vertical neighbors $i,j$, there are exactly $2m$ boundary indices $\alpha$ s.t. $\left(\mathbf{D}^{(v)}_{\alpha,i}\right)^{\rm T}\mathbf{D}^{(v)}_{\alpha,j}\neq 0$, and $\left\|\left(\mathbf{D}^{(v)}_{\alpha,i}\right)^{\rm T}\mathbf{D}^{(v)}_{\alpha,j}\right\|_{2}\leq 2$. Moreover,
$\left\|\left((\mathbf{D}_{\alpha}^{(v)}\mathbf{x})\mathbf{D}_{\alpha,i}^{(v)}-(\mathbf{D}_{\alpha}^{(h)}\mathbf{x})\mathbf{D}_{\alpha,i}^{(h)}\right)^{\rm T}\left((\mathbf{D}_{\alpha}^{(v)}\mathbf{x})\mathbf{D}_{\alpha,j}^{(v)}-(\mathbf{D}_{\alpha}^{(h)}\mathbf{x})\mathbf{D}_{\alpha,j}^{(h)}\right)\right\|_{2}$
$\leq\left(\sqrt{2}|\mathbf{D}_{\alpha}^{(v)}\mathbf{x}|+\sqrt{2}|\mathbf{D}_{\alpha}^{(h)}\mathbf{x}|\right)^{2}\leq 4\left((\mathbf{D}_{\alpha}^{(v)}\mathbf{x})^{2}+(\mathbf{D}_{\alpha}^{(h)}\mathbf{x})^{2}\right)$
Therefore, recall Eq. 36, it holds that
(38) $\begin{split}\left\|H_{ij}(\mathbf{x})\right\|_{2}\leq~{}&\delta\sum_{\alpha\text{ in boundary}}\left((\mathbf{D}_{\alpha}^{(v)}\mathbf{x})^{2}+(\mathbf{D}_{\alpha}^{(h)}\mathbf{x})^{2}+\varepsilon\right)^{-3/2}\\ &\cdot\left[4\left((\mathbf{D}_{\alpha}^{(v)}\mathbf{x})^{2}+(\mathbf{D}_{\alpha}^{(h)}\mathbf{x})^{2}\right)+4\varepsilon\right]\\ \leq~{}&4\delta\sum_{\alpha\text{ in boundary}}\left((\mathbf{D}_{\alpha}^{(v)}\mathbf{x})^{2}+(\mathbf{D}_{\alpha}^{(h)}\mathbf{x})^{2}+\varepsilon\right)^{-1/2}\leq 8m\delta\varepsilon^{-1/2}.\end{split}$
Similar result holds when $i,j$ are horizontal neighbors. Also notice $H_{jk}(\mathbf{x})=0$ when $j\not\sim k$, where we denote $j\sim k$ if $j\neq k$ are neighboring blocks. Now substitute the above inequalities into Eq. 37, we have
(39) $\begin{split}&-\left(\nabla_{j}u(\mathbf{x})\right)^{\rm T}\nabla_{j}\phi(\mathbf{x}_{i})+\left(\nabla_{j}u(\mathbf{x})\right)^{\rm T}\mathcal{L}(\nabla_{j}u)(\mathbf{x})\\ \geq~{}&\lambda M_{jj}\|\nabla_{j}u(\mathbf{x})\|_{2}^{2}-\sum_{k\neq j}(\lambda M_{jk}+8m\delta\varepsilon^{-1/2}{\bf{1}}_{j\sim k})\|\nabla_{j}u(\mathbf{x})\|_{2}\|\nabla_{k}u(\mathbf{x})\|_{2}.\end{split}$
Consider $\mathbf{x}$ where $\|\nabla_{j}u(\mathbf{x})\|_{2}$ reaches its maximum, i.e. $\|\nabla_{j}u(\mathbf{x})\|_{2}=\left\|\nabla_{j}u\right\|_{L^{\infty}}$. The first order optimality condition reads
$0=\nabla\left(\|\nabla_{j}u(\mathbf{x})\|_{2}^{2}\right)=\nabla\nabla_{j}u(\mathbf{x})\cdot\nabla_{j}u(\mathbf{x}),$
and the second order optimality condition reads
$0\geq\Delta\left(\|\nabla_{j}u(\mathbf{x})\|_{2}^{2}\right)=2\|\nabla\nabla_{j}u(\mathbf{x})\|_{\rm F}^{2}+2(\nabla_{j}u(\mathbf{x}))^{\rm T}\Delta\nabla_{j}u(\mathbf{x}).$
So that $(\nabla_{j}u(\mathbf{x}))^{\rm T}\Delta\nabla_{j}u(\mathbf{x})\leq 0$ since $\|\nabla\nabla_{j}u(\mathbf{x})\|_{\rm F}^{2}\geq 0$. Under these conditions,
So that at the maximum point, Eq. 39 reads
$\lambda M_{jj}\|\nabla_{j}u\|_{L^{\infty}}^{2}-$ $\sum_{k\neq j}(\lambda M_{jk}+8m\delta\varepsilon^{-1/2}{\bf{1}}_{j\sim k})\left\|\nabla_{j}u\right\|_{L^{\infty}}\|\nabla_{k}u(\mathbf{x})\|_{2}$